# ContextFormer: Redefining Efficiency in Semantic Segmentation

### v1

@@ Abstract
Semantic segmentation assigns labels to pixels in images, a critical yet challenging task in computer vision. Convolutional methods, although capturing local dependencies well, struggle with long-range relationships. Vision Transformers (ViTs) excel in global context capture but are hindered by high computational demands, especially for high-resolution inputs. Most research optimizes the encoder architecture, leaving the bottleneck underexplored—a key area for enhancing performance and efficiency. We propose ContextFormer, a hybrid framework leveraging the strengths of CNNs and ViTs in the bottleneck to balance efficiency, accuracy, and robustness for real-time semantic segmentation. The framework’s efficiency is driven by three synergistic modules: the Token Pyramid Extraction Module (TPEM) for hierarchical multi-scale representation, the Transformer and Modulating DepthwiseConv (Trans-MDC) block for dynamic scale-aware feature modeling, and the Feature Merging Module (FMM) for robust integration with enhanced spatial and contextual consistency. Extensive experiments on ADE20K, Pascal Context, CityScapes, and COCO-Stuff datasets show ContextFormer significantly outperforms existing models, achieving state-of-the-art mIoU scores, setting a new benchmark for efficiency and performance. The codes will be made publicly available upon acceptance.
@@ 1 Introduction
Semantic segmentation, a key task in computer vision, assigns each pixel in an image a specific semantic label [45]. Traditional approaches [25, 24, 51, 5, 1] address this challenge by employing stacks of local convolutional kernels to capture long-range spatial dependencies. However, with the advent of Vision Transformers (ViTs), semantic segmentation has undergone a significant transformation [18]. These Transformer-based approaches [7, 62] excel in capturing global context, showcasing superior performance compared to traditional methods. However, their high computational and memory demands pose significant challenges, particularly for real-world applications involving high-resolution input data [11, 39, 36, 42], making them less practical for deployment at scale.
To address this, several efficient strategies, such as local/window-based attention [53], axial attention [26] and lightweight attention mechanisms [43, 37, 36, 48, 69] have been developed. These methods aim to reduce the computational load, but they still struggle to meet the low-latency and the segmentation accuracy essential for real-world applications, especially on high-resolution images (as demonstrated in Fig. 1). To mitigate the computational costs associated with high-resolution processing in a transformer based approach, the authors in [69] applied global attention at a down-scaled resolution of 1///64 of the original input, a compromise that inevitably degrades segmentation performance. Conversely, several proposed lightweight Convolutional Neural Networks (CNNs) based models [14, 15, 41, 70, 27] benefit via leveraging asymmetric or depth-wise separable convolutions [52, 28] with linear complexity relative to the input size, making them better suited for edge device deployment. However, these CNN-based approaches rely on local receptive fields for feature extraction, which constrains their segmentation accuracy.
To further improve the model performance, the researchers explored hybrid transformer-CNN architectures, leveraging the complementary strengths of both paradigms. In TrSeg [30], ResNet serves as the backbone network, while transformer layers are incorporated within the decoder to merge contextual information specific to traffic scenes. Wang et al. [61] employed the Swin Transformer [39] to transform an image into a sequence of feature embeddings, utilizing a CNN decoder to reconstruct the spatial dimensions of the feature maps. Several approaches [20, 6, 42, 69] combined them by using CNNs as encoder/backbone for feature extraction and ViTs as bottleneck for improving the efficiency. They had shown improvement in the performance, but at the cost of increased latency. However, despite efforts to balance accuracy, speed, and computational cost, the potential of combining CNNs and ViTs within the bottleneck to synergize long-range and short-range spatial context modeling — offering both performance and efficiency—remains underexplored.
Building on the limitations of prior approaches, we present ContextFormer, a Transformer-based framework for semantic segmentation that efficiently captures both local and global context in the bottleneck design, achieving a balanced trade-off between efficiency, accuracy, and robustness. By integrating three key components—the Token Pyramid Extraction Module (TPEM), Transformer and Modulating DepthwiseConv (Trans-MDC) block, and Feature Merging Module (FMM)—the framework ensures efficient real-time segmentation of high-resolution imagery, delivering high-fidelity predictions tailored for real-world deployment. Specifically, TPEM constructs a hierarchical feature pyramid by fusing image data, gradient magnitude maps, and edge features, leveraging cascaded MobileNetV2 blocks [52] for efficient multi-scale feature extraction with minimal overhead. The Trans-MDC block processes multi-scale tokens from TPEM using a hybrid architecture to dynamically align global dependencies with fine-grained details, producing scale-aware, contextualized features. Complementing this, the FMM establishes robust feature hierarchies through adaptive channel redistribution, enhancing spatial consistency and representational capacity via a gated-shifting mechanism. These enriched features enable ContextFormer to achieve high-fidelity predictions with computational efficiency, optimized for real-world applications.
@@
To rigorously evaluate the efficacy of our proposed approach, we conduct a series of experiments on complex segmentation benchmarks, including ADE20K [73], Pascal Context [46], and CityScapes [9]. To further validate the generalization of our model, we also apply it to the COCO-Stuff [2] dataset. The primary contributions of the work can be summarized as follows:
@@
  • We introduce a novel bottleneck design utilizing Trans-MDC blocks, which dynamically capture global dependencies and scale-aware semantics. These blocks, coupled with the TPEM for efficient multi-scale feature extraction and the FMM for adaptive feature integration, ensure robust hierarchical representations essential for dense prediction tasks.
@@ 3 Proposed Method
This section presents ContextFormer, an efficient, robust, and powerful segmentation framework that avoids computationally intensive components. As shown in Fig. 2, it is composed of three key modules: (1) a Token Pyramid Encoding Module that captures high-resolution coarse features and low-resolution fine features and generates a feature pyramid, (2) a hybrid Trans-MDC block that efficiently generates scale-aware semantics, and (3) a lightweight Feature Merging Module that fuses these multi-level local and global semantic features to produce semantic segmentation mask via shifted gated mechanism. Next, the details of the model parts are explained below.  [^1] For better readability and ease, non-linear activation functions, batch normalization and dropout layers are not mentioned in the equations and figures.
3.1 Token Pyramid Encoding Module
@@
Considering the efficiency constraints, the proposed Token Pyramid Encoding module (TPEM) incorporates a series of stacked MobileNet blocks [52], specifically inverted residual blocks. However, unlike MobileNets, the proposed TPEM employs fewer blocks to establish a lightweight hierarchical feature representation. As demonstrated in Fig. 2, an input image ${X\in\mathbb{R}^{5\times H\times W}}$ consists of stacked 3 channel (RGB) image along with the gradient magnitude and edge maps to enhance the semantic richness, where $H$ and $W$ are the height and width of the image, respectively. The TPEM module initiates by passing this concatenated feature map via a series of MobileNetV2 blocks resulting in multi-scale features ${\{S_{1},S_{2},...,S_{L}\}}$ at different levels, where the number of levels is denoted by $L$ and the hierarchical feature at each level $L$ denoted as $\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times C_{i}$, where $i\in\{1,2,3,4\}$. These multi-scale features encompass high-resolution coarse representations alongside low-resolution fine-grained details, which collectively enhance the efficacy of semantic segmentation. To further minimize the computational overhead, an average pooling operation is applied to downsample the extracted tokens to a significantly reduced size, for instance, to $\frac{1}{64\times 64}$ of the original input size. Finally, the same sized tokens from different scales are concatenated along the channel dimension to enhance the inter-channel dependencies and produce refined tokens for subsequent blocks. The overall operation of the TPEM is defined as:
$X_{f}=<S_{1}^{\varphi},S_{2}^{\varphi},S_{3}^{\varphi},S_{4}^{\varphi}>$ (1)
Where, $\varphi$ denotes the average pooling operation applied to the corresponding feature map, and <.><.>< . > represents the concatenation operation. This combined local feature set is then passed to the Trans-MDC block to extract global and enhanced multi-scale semantics.
3.2 Trans-MDC Block
The average pooled tokens from various scales are unified to a consistent resolution and thereafter concatenated to serve as the input for the Vision Transformer (ViTs) and the Modulating DepthwiseConv block (Trans-MDC). This enables the network to capture a full-image receptive field and extract the rich semantic global and local information in an efficient way. As illustrated in Fig. 2, Trans-MDC block takes as input the combined feature set output of TPEM ($X_{f}$) and modulates the mutli-head attention output of the ViT. Basically, the MDC block draws inspiration from the architectural framework of ViTs, while being composed exclusively of convolutional layers: a $3\times 3$ depthwise convolution, a $1\times 1$ depthwise convolution, and a $3\times 3$ depthwise separable convolution. This design enables the model to learn rich and nuanced features essential for effective semantic segmentation. The inclusion of depthwise convolutions in our design is pivotal due to their significant reduction in computational cost and memory usage. Here, each convolution operation contributes distinct advantages: the $3\times 3$ depthwise convolution excels at capturing spatial semantics, while the $1\times 1$ depth-wise convolution facilitates the learning of depth-wise cross-channel information. Additionally, the $3\times 3$ depth-wise separable convolution integrates both the spatial and channel-wise semantics. Thereafter, the combined semantics, $\delta_{c}^{{}^{\prime}}$ obtained via aggregating the information from all the three layers are subsequently passed through a channel attention module as shown in Eq. 3, which further enhances the global contextual information, thereby boosting the robustness of the extracted features. The overall operation of the MDC block is defined as:
@@ 3.2 Trans-MDC Block
$X_{MDC}=\Gamma_{2}(\Gamma_{1}(\varphi_{g}(\delta_{c}^{{}^{\prime}})))\odot\delta_{c}^{{}^{\prime}}$ (3)
@@
Further, as demonstrated in Fig. 2, the ViT block also ingests the pooled feature maps ($X_{f}$), whereby the global self-attention mechanism facilitates the exchange of information among tokens across the spatial dimension. To maintain the spatial integrity of the tokens and to minimize the frequency of reshaping operations typically associated with ViT architectures, we substitute the conventional linear layers with 1×1$\times$1 ×1 convolutional layers. In our implementation of multi-head attention within the ViT framework, we configure the dimensions for the $Q,K,V$ representations to be 16, 16, and 32, respectively. This dimensional reduction serves to further decrease the overall computational complexity of the model. Additionally, we opt for batch normalization over layer normalization for faster inference. It is important to note that batch normalization is applied following each convolutional layer throughout the network, with the exception of the final output convolution layer. Furthermore, we replace the GELU non-linear activation function originally employed in ViT with ReLU6 [28] in our architecture, thereby enhancing performance while retaining effective activation characteristics.
@@
Where, $Attention=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$. Finally, the output features from both the blocks are combined to achieve efficient and enhanced features.
$X_{f}^{{}^{\prime}}=X_{ViT}+X_{MDC}$ (5)
The unified output semantics from the Trans-MDC block are thereby passed through the feed-forward network. For the feed-forward network, following [29, 65], we have integrated depth-wise convolution layer between $1\times 1$ convolution layers and to further minimize the computational complexity, expansion factor of two is incorporated which helps the Trans-MDC block to augment the capturing of global semantics. The architectural details about FFN are given in the supplementary. The overall operation of Trans-MDC block can be summarized as follows:
@@ 3.3 Feature Merging Module and Segmentation Head
After obtaining robust and enhanced semantics from the Trans-MDC block, Feature Merging Module (FMM) is incorporated to merge the complementary features from the Token Pyramid Encoding Module (TPEM) and Trans-MDC block. FMM module tries to alleviate the semantic gap of the different incoming information and allows the model to have both the local and global context information via its gated shifted mechanism. To this end, in FMM, firstly the local features from TPEM are passed through a $1\times 1$ convolution layer. Further, to get a weight matrix for the global features, a $1\times 1$ convolution layer followed by Sigmoid activation function is applied on the global features/semantics from Trans-MDC block. Subsequently, the global semantics are integrated into the local tokens via the Hadamard product, and these global semantics are further summed, as shown in Fig. 2, with the local features via a $1\times 1$ convolution layer to focus on the relevant semantics. The overall operation of FMM module is summarized as follows:
@@
Where, the first term before the Hadamard product denotes the processing on the local tokens incoming from the TPEM module, the second term denotes the gated operation being performed on the features from the Trans-MDC block to obtain a weight matrix, and $\sigma$ denotes Sigmoid activation function.
@@ 4.1 Datasets and Measures
Following the recent literature [31, 3, 69] we report the results using the standard common measures: Mean Intersection over Union (mIoU) for segmentation accuracy, Giga Floating Point Operations per Second (GFLOPs), latency and number of parameters.
@@ 4.2 Implementation Details
Our implementation is based on PyTorch and MMSegmentation toolbox [8]. All the models are first pretrained on ImageNet-1K dataset [12], including our proposed ContextFormer [^2] On ImageNet-1K dataset, ContextFormer achieves 66.11% Top-1 accuracy. Details are provided in the supplementary material. model. Further, they are fined-tuned on semantic segmentation datasets. Batch-Normalization layers are used after almost each convolution layer except the last output layer. For the ADE20K dataset, we follow data augmentations same as in [62] for fair comparison. Moreover, for ADE20K dataset, we use batch size 16, and 160K scheduler by following [62] and [69]. Various augmentations have been applied such as random scaling, random cropping, random horizontal flip, random resize etc. For the CityScapes dataset, same data augmentations are followed as in [62, 69]. Images are resized and rescaled with same crop size i.e. $1024\times 512$. It should be noted that for all datasets and models, we set $1.2\times 10^{-4}$ as the initial learning rate with the weight decay set to $0.01$. However, initial learning rate for the CityScapes dataset is set to $3\times 10^{-4}$. For the PASCAL Context and COCO-Stuff datasets, 80K training iterations are implemented. Additionally, for both datasets, same data augmentations and training settings are incorporated as in [8] and the training images are resized and cropped to $512\times 512$ and $480\times 480$ for COCO-Stuff and PASCAL Context datasets, respectively.
Table 1: Quantitative results of various models on ADE20K validation set [73]. * indicates results from models trained with $448\times 448$ input size. For fair comparison with SegFormer and TopFormer, batch size of 16 is considered for ContextFormer. Whereas, batch size of 32 is considered for CNN based models. GFLOPs are reported for input resolution of $512\times 512$. Single-scale inference is used to report mIoU. ‘-’ means the unreported results, and ContextFormer(GME) is proposed model with concatenated edge and magnitude maps.
@@
ContextFormer (Ours) TPEM 34.6 0.6 1.7M 16
@@
ContextFormer(GME) (Ours) TPEM 35.0 0.6 1.7M 16

### v2

@@ Abstract
Semantic segmentation assigns labels to pixels in images, a critical yet challenging task in computer vision. Convolutional methods, although capturing local dependencies well, struggle with long-range relationships. Vision Transformers (ViTs) excel in global context capture but are hindered by high computational demands, especially for high-resolution inputs. Most research optimizes the encoder architecture, leaving the bottleneck underexplored—a key area for enhancing performance and efficiency. We propose ContextFormer, a hybrid framework leveraging the strengths of CNNs and ViTs in the bottleneck to balance efficiency, accuracy, and robustness for real-time semantic segmentation. The framework’s efficiency is driven by three synergistic modules: the Token Pyramid Extraction Module (TPEM) for hierarchical multi-scale representation, the Transformer and Branched DepthwiseConv (Trans-BDC) block for dynamic scale-aware feature modeling, and the Feature Merging Module (FMM) for robust integration with enhanced spatial and contextual consistency. Extensive experiments on ADE20K, Pascal Context, CityScapes, and COCO-Stuff datasets show ContextFormer significantly outperforms existing models, achieving state-of-the-art mIoU scores, setting a new benchmark for efficiency and performance. The codes will be made publicly available upon acceptance.
@@ 1 Introduction
Semantic segmentation is a fundamental task in computer vision that involves dense pixel-wise classification [46]. Traditional methods [25, 24, 52, 5, 1] rely on Convolutional Neural Networks (CNNs) to extract hierarchical feature representations using local receptive fields. However, convolutional operators inherently struggle with capturing long-range dependencies, leading to suboptimal segmentation accuracy in complex scenes. With the advent of Vision Transformers (ViTs) [18], segmentation models have demonstrated superior global context modeling [7, 63]. Despite their effectiveness, ViTs exhibit quadratic complexity in self-attention computations, leading to excessive memory and compute requirements for high-resolution inputs [11, 40, 35, 43], making real-time deployment challenging.
To address these computational constraints, researchers have explored efficient transformer variants such as local/window-based attention [54], axial attention [26], and lightweight attention mechanisms [44, 36, 35, 49, 69]. These methods partially alleviate complexity but still suffer from degradation in segmentation accuracy due to constrained receptive fields and reduced feature expressiveness, especially on high-resolution images (as demonstrated in Fig. 1). To mitigate the computational costs associated with high-resolution processing in a transformer based approach, TopFormer [69] applies self-attention at a 1/64 downscaled resolution, reducing computational overhead at the cost of losing fine-grained spatial details. On the other hand, mobile-friendly CNN-based models  [14, 15, 42, 70, 27] leverage depth-wise separable convolutions  [53, 28] for linear complexity scaling, but their local feature aggregation limits long-range contextual reasoning, ultimately reducing segmentation quality.
To bridge this gap, hybrid CNN-Transformer architectures have been proposed. Some methods integrate CNN backbones with Transformer decoders [29, 62], while others employ Transformer bottlenecks for global information processing [20, 6, 43, 69]. While these approaches improve efficiency and expressiveness, they often struggle with an inherent trade-off: Transformer-based bottlenecks can introduce redundancy in global attention computation, while CNN-based bottlenecks, limited by local receptive fields, may fail to capture long-range dependencies effectively thus limiting their ability to simultaneously optimize efficiency, accuracy, and scalability in real-world segmentation tasks.
Building on these insights, we propose ContextFormer, a bottleneck-optimized segmentation framework that efficiently integrates global and local feature processing within a lightweight computational design. It comprises three key components: the Token Pyramid Extraction Module (TPEM), Transformer and Branched Depthwise Conv (Trans-BDC) block, and Feature Merging Module (FMM). TPEM constructs multi-scale token representations by fusing raw image intensity, gradient magnitude, and edge features via cascaded MobileNetV2 blocks [53], ensuring efficient feature encoding. The Trans-BDC block enhances contextual reasoning while preserving computational efficiency by integrating modulated branched depthwise convolutions with lightweight global attention. FMM adaptively redistributes spatial and channel-wise features across resolutions, refining hierarchical representations for robust segmentation. By synergizing these components within the bottleneck design, ContextFormer achieves superior efficiency-accuracy trade-offs, addressing key limitations in hybrid CNN-Transformer architectures.
@@
To rigorously evaluate the efficacy of our proposed approach, we conduct a series of experiments on complex segmentation benchmarks, including ADE20K [73], Pascal Context [47], CityScapes [9], and COCO-Stuff [2]. To further highlight the generalizability of our model, we validate its performance by conducting object detection experiments on the COCO [37] dataset. The primary contributions of the work can be summarized as follows:
@@
  • We introduce a novel bottleneck design utilizing Trans-BDC blocks, which dynamically capture global dependencies and scale-aware semantics. These blocks, coupled with the TPEM for efficient multi-scale feature extraction and the FMM for adaptive feature integration, ensure robust hierarchical representations essential for dense prediction tasks.
@@ 3 Proposed Method
This section presents ContextFormer, an efficient, robust, and powerful segmentation framework that avoids computationally intensive components. As shown in Fig. 2, it is composed of three key modules: (1) a Token Pyramid Extraction Module that captures high-resolution coarse features and low-resolution fine features and generates a feature pyramid, (2) a hybrid Trans-BDC block that efficiently generates scale-aware semantics, and (3) a lightweight Feature Merging Module that fuses these multi-level local and global semantic features to produce semantic segmentation mask via shifted gated mechanism. Next, the details of the model parts are explained below. [^1] For better readability and ease, non-linear activation functions, batch normalization and dropout layers are not mentioned in the equations and figures.
3.1 Token Pyramid Extraction Module
@@
Considering the efficiency constraints, the proposed Token Pyramid Extraction Module (TPEM) incorporates a series of stacked MobileNet blocks [53], specifically inverted residual blocks. However, unlike MobileNets, the proposed TPEM employs fewer blocks to establish a lightweight hierarchical feature representation. As demonstrated in Fig. 2, an input image ${X\in\mathbb{R}^{5\times H\times W}}$ consists of stacked 3 channel (RGB) image along with the gradient magnitude and edge maps to enhance the semantic richness, where $H$ and $W$ are the height and width of the image, respectively. The TPEM module initiates by passing this concatenated feature map via a series of MobileNetV2 blocks resulting in multi-scale features ${\{S_{1},S_{2},...,S_{L}\}}$ at different levels, where the number of levels is denoted by $L$ and the hierarchical feature at each level $L$ denoted as $\frac{H}{2^{i+1}}\times\frac{W}{2^{i+1}}\times C_{i}$, where $i\in\{1,2,3,4\}$.
These multi-scale features encompass high-resolution coarse representations alongside low-resolution fine-grained details, which collectively enhance the efficacy of semantic segmentation. To further minimize the computational overhead, an average pooling operation is applied to downsample the extracted tokens to a significantly reduced size, for instance, to $\frac{1}{64\times 64}$ of the original input size. Finally, the same sized tokens from different scales are concatenated along the channel dimension to enhance the inter-channel dependencies and produce refined tokens for subsequent blocks. The overall operation of the TPEM is defined as:
$X_{f}=\langle S_{1}^{\varphi},S_{2}^{\varphi},S_{3}^{\varphi},S_{4}^{\varphi}\rangle$ (1)
[…]
3.2 Trans-BDC Block
The Trans-BDC block serves as an efficient bottleneck, integrating branched depthwise convolutions with lightweight self-attention to balance local and global feature extraction while reducing computational cost. As illustrated in  Fig. 2, to enhance efficiency, multi-scale features from the Token Pyramid Extraction Module (TPEM) undergo average pooling, ensuring a unified resolution before being processed in two parallel branches: a Branched Depthwise Convolution Module (BDC) for local feature extraction and a lightweight Vision Transformer (ViT) for global self-attention. Pooling at this stage reduces spatial redundancy, lowering FLOPs before deeper processing.
Basically, the BDC module employs three parallel depthwise convolutions: a $3\times 3$ depthwise convolution for spatial semantics, a $1\times 1$ depthwise convolution for cross-channel interactions, and a $3\times 3$ depthwise separable convolution for efficient feature fusion. These outputs are aggregated and refined via a channel attention mechanism, enhancing contextual understanding with minimal computational overhead and the overall operation of BDC is defined as:
@@ 3.2 Trans-BDC Block
$X_{BDC}=\Gamma_{2}(\Gamma_{1}(\varphi_{g}(\delta_{c}^{{}^{\prime}})))\odot\delta_{c}^{{}^{\prime}}$ (3)
@@
Simultaneously, the ViT branch applies self-attention on the pooled features, but with key optimizations: $1\times 1$ convolutions replace MLPs, reducing reshaping operations; lower-dimensional projections ($Q,K,V$ set to 16, 16, and 32, respectively) decrease memory footprint; batch normalization substitutes layer normalization for faster inference; and GELU activation is replaced with ReLU6  [28], further improving efficiency.
@@
Where, $Attention=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$. The outputs from the BDC and ViT branches are fused to form $X_{f}^{{}^{\prime}}$, forming a refined feature representation that is then processed by a Feed-Forward Network (FFN) enhanced with depthwise convolutions and the overall operation of this block can be summarised as:
@@
This reduces computational complexity while maintaining strong representational power. By leveraging branched feature extraction, token pooling, and optimized self-attention, Trans-BDC significantly lowers the computational burden in the bottleneck compared to conventional CNN or transformer-based architectures while preserving segmentation accuracy.
@@ 3.3 Feature Merging Module and Segmentation Head
After obtaining robust and enhanced semantics from the Trans-BDC block, Feature Merging Module (FMM) is incorporated to merge the complementary features from the Token Pyramid Extraction Module (TPEM) and Trans-BDC block. FMM module tries to alleviate the semantic gap of the different incoming information and allows the model to have both the local and global context information via its gated shifted mechanism. To this end, in FMM, firstly the local features from TPEM are passed through a $1\times 1$ convolution layer. Further, to get a weight matrix for the global features, a $1\times 1$ convolution layer followed by Sigmoid activation function is applied on the global features/semantics from Trans-BDC block. Subsequently, the global semantics are integrated into the local tokens via the Hadamard product, and these global semantics are further summed, as shown in Fig. 2, with the local features via a $1\times 1$ convolution layer to focus on the relevant semantics. The overall operation of FMM module is summarized as follows:
@@
Where, the first term before the Hadamard product denotes the processing on the local tokens incoming from the TPEM module, the second term denotes the gated operation being performed on the features from the Trans-BDC block to obtain a weight matrix, and $\sigma$ denotes Sigmoid activation function.
@@ 4.1 Datasets and Measures
Following the recent literature [30, 3, 69] we report the results using the standard common measures: Mean Intersection over Union (mIoU) for segmentation accuracy, Giga Floating Point Operations per Second (GFLOPs), latency and number of parameters. Details for object detection task are provided in Sec. 4.7.
@@ 4.2 Implementation Details
Our implementation is based on PyTorch and MMSegmentation toolbox [8]. All the models are first pretrained on ImageNet-1K dataset [12], including our proposed ContextFormer [^2] On ImageNet-1K dataset, ContextFormer achieves 66.4% Top-1 accuracy. Details are provided in the supplementary material. model. Further, they are fined-tuned on semantic segmentation datasets. Batch-Normalization layers are used after almost each convolution layer except the last output layer. For the ADE20K dataset, we follow data augmentations same as in [63] for fair comparison. Moreover, for ADE20K dataset, we use batch size 16, and 160K scheduler by following [63] and [69]. Various augmentations have been applied such as random scaling, random cropping, random horizontal flip, random resize, etc. For the CityScapes dataset, the same data augmentations are followed as in [63, 69]. Images are resized and rescaled with same crop size i.e. $1024\times 512$. It should be noted that for all datasets and models, we set $1.2\times 10^{-4}$ as the initial learning rate with the weight decay set to $0.01$. However, the initial learning rate for the CityScapes dataset is set to $3\times 10^{-4}$. For the PASCAL Context and COCO-Stuff datasets, 80K training iterations are implemented. Additionally, for both datasets, the same data augmentations and training settings are incorporated as in [8] and the training images are resized and cropped to $512\times 512$ and $480\times 480$ for COCO-Stuff and PASCAL Context datasets, respectively.
Table 1: Quantitative results of various models on ADE20K validation set [73]. * indicates results from models trained with $448\times 448$ input size. For fair comparison with SegFormer, TopFormer and SeaFormer batch size of 16 is considered for ContextFormer. Whereas, batch size of 32 is considered for CNN based models. GFLOPs are reported for input resolution of $512\times 512$. ‘-’ means the unreported results. Single-scale inference is used to report mIoU.
@@
ContextFormer (Ours) TPEM 34.8 0.6 1.7M 15
@@
ContextFormer(GME) (Ours) TPEM 35.0 0.6 1.7M 15